# Neural Diffusion Processes for Physically Interpretable Survival Prediction

### v2

@@ Abstract
We compare DeepFHT with Cox survival model using synthetic and real-world datasets. The method achieves predictive accuracy on par with state-of-the-art approaches, while maintaining a physics-based interpretable parameterization that elucidates the relation between input features and risk. This combination of stochastic process theory and deep learning provides a principled avenue for modeling survival phenomena in complex systems.
@@ I Introduction
Survival analysis is central in many applications across medicine, engineering, economics and finance. It concerns time-to-event modeling: given a process that can generate an event of interest (e.g., death from disease, failure due to wear), the goal is to estimate the probability that an event occurs at time $t$ for an individual described by input variables (or features, or covariates). Unlike standard regression settings, survival data are characterized by censoring, which means that for some instances, the exact event time is not observed (for example, when they remain event-free at the end of the study), and only the last recorded follow-up time is available.
Traditional approaches to survival modeling rely on strong statistical assumptions linking input variables and risk. The Cox proportional hazards (CoxPH) model [1] remains the most widely used and best established method. The proportional hazards assumption implies that the instantaneous risk of an event for two individuals differs by a constant factor over time. The Cox model is also linear, making it clear how each single input variable affects the outcome, but at the expense of missing interactions between features. In its original form, this relation is modeled through a linear regression on the features, though many extensions have been developed to relax linearity and improve performance in high-dimensional settings [2, 3, 4].
@@
Among these, first hitting time (FHT) models are of particular interest for this work. FHT models assume that event times are distributed as the times of first passage of a latent stochastic process, underlying the event-generating process observed, through a barrier. This approach overcomes the PH assumption, since FHT distributions for many stochastic processes allow for time-varying hazard ratios between different individuals (e.g., Lévy distribution, inverse Gaussian distribution).
@@
A branch of FHT methods aims to model the distribution of first hitting times directly, using machine learning techniques [10] and deep neural networks [11] to achieve strong predictive performance, but often failing to give interpretable insights into the nature of the underlying process and how features relate to it, due to the black box nature of most deep learning frameworks. Conversely, parametric FHT alternatives [12, 13, 14] currently suffer from the opposite drawbacks. These models learn the process parameters on which the distribution of first hitting times depends from data, using simpler regression and machine learning techniques, yielding more physically interpretable results at the expense of model expressiveness.
The present work proposes a framework that combines these complementary strengths. Our model, DeepFHT, integrates parametric FHT modeling with deep neural networks, which are used to map input features to the parameters of the underlying stochastic process. To our knowledge, this is the first application of modern deep learning techniques to parametric FHT models for survival prediction. The hybrid design leverages the representational capacity of neural networks while preserving interpretability through the physical meaning of process parameters. In this way, DeepFHT offers a controlled alternative to black-box survival models: while it does not eliminate the challenges of interpretability, it grounds predictions in a dynamical description that connects input features to the evolution of an underlying stochastic process.
@@
We demonstrate the capability of our framework on synthetic and publicly available clinical survival datasets, comparing performance against the Cox model. DeepFHT achieves competitive accuracy across clinical benchmarks and clearly outperforms Cox regression on synthetic data designed to probe the model’s capabilities in scenarios with nonlinearities and non-proportional hazards. Finally, we assess interpretability by examining correlations between predicted process parameters, empirical risk and features. The results show that similarity between patient outcomes is intuitively encoded as distance in the space of parameters, and that DeepFHT can recover clinically meaningful relationships and has the potential to highlight new correlations in complex, high-dimensional data.
@@
The remainder of the paper is organized as follows. In Section II and Section III we briefly present the central problem of survival analysis and the necessary theory of stochastic processes. Section IV presents the proposed model and the experimental setup, while results and the notion of physics-based interpretability of the model are discussed in Section V.
@@ II Survival analysis framework
Survival data are typically represented as triplets $(\mathbf{x}_{i},\delta_{i},T_{i})$ for $i=1,\dots,n$, where:
  • $\mathbf{x}_{i}\in\mathbb{R}^{m}$ is a vector of features
  • $\delta_{i}\in\{0,1\}$ is an indicator that denotes whether the event was observed ($\delta_{i}=1$) or the instance was right censored, i.e., end of observation without occurrence of the event ($\delta_{i}=0$)
  • $T_{i}$ is the time of event occurrence or censoring
  • $n$ is the dataset size
The goal of survival analysis is to predict the probability distribution of event times $f(t)$ and some relevant, related quantities:
@@
The survival function describes the probability of surviving beyond time $t$, while the hazard captures the instantaneous risk of event occurrence given survival up to $t$. Estimation typically focuses on conditional forms $S(t|\mathbf{X})$ and $h(t|\mathbf{X})$, which quantify the dependence of survival dynamics on features.
@@ III First hitting time models for survival analysis
A natural probabilistic representation of survival processes is provided by the theory of first hitting times. Consider a 1D continuous-time stochastic process $\{X(t)\}_{t\geq 0}$, with $X(t)\in\mathbb{R}$, described by a probability distribution $p(x,t)$. Let $\Gamma:\{x\leq x_{b}\}\subset\mathbb{R}$ be a target set and $X(t=0)=x_{0}>x_{b}$ the initial condition. In this setup, $x_{b}$ can be considered as a barrier between the target set and the initial condition and this is how it will be referred to in the rest of the paper. The first hitting time is then
@@
that is, the earliest time at which the process crosses the barrier in $x_{b}$. When the barrier is considered to be absorbing (i.e., killing the process upon transition), the distribution $f(T)$ of first hitting times coincides with the event time distribution in survival analysis. This condition is easily imposed by the Dirichlet condition $p(x_{b},t\mid x_{0})=0$. Without losing generalizability, we will always consider positive initial conditions $x_{0}>0$ and an absorbing barrier in $x_{b}=0$. This amounts to a generic choice for a process in $X(t)\in\mathbb{R}$.
Since all relevant survival functions can be derived from the distribution of first hitting times $f(T)$, it is crucial to have analytical solutions for this quantity. In this work, we consider two stochastic processes whose theory is well developed and for which an analytical solution is available: the Brownian motion and the arithmetic Brownian motion (Brownian motion with drift). For a more complete and detailed discussion of FHT theory and models, including but not limited to their application to survival analysis, we refer to [15, 12, 16].
@@ III.1 First hitting time and survival distributions for Brownian motion
    the distribution of first hitting times (i.e., event times)
@@ III.2 First hitting time and survival distributions for arithmetic Brownian motion
with constant diffusion coefficient $D>0$. For an initial condition $p(x,0\mid x_{0})=\delta(x-x_{0})$, the solution is Gaussian with the mean shifted linearly in time:
@@
Introducing an absorbing barrier at $x_{b}=0$ modifies the solution which becomes, via the method of images:
@@
    where $\Phi(\cdot)$ denotes the standard normal cdf
@@
    which corresponds to an inverse Gaussian distribution
@@ IV.1 DeepFHT
The dimension $d$ of the output vector $\mathbf{p}$ depends on the choice of parametrization of the FHT distribution. As anticipated in Section III, in this paper we will focus on Brownian and arithmetic Brownian motion, identified by their FHT distributions 5,10, with the following parameterizations:
@@
Notably, both parameterizations above allow for non-proportional hazards, reflecting the flexibility of FHT-based models to capture complex time dependencies in risk. Figure 1 illustrates a representative model output, where survival functions are computed from the estimated distributional parameters for a set of individuals.
It is also worth noting that this model can easily be extended to any other probability distribution of choice. It is sufficient to specify the functional form and its parameterization and to modify the size of the output layer accordingly. In Sec. IV.3 we take advantage of this flexibility to compare results obtained with Lévy and Inverse Gaussian distributions against those obtained with failure density distributions that have no direct interpretation in the first hitting time framework.
@@ IV.2 Loss function: Brier loss
Training is performed by minimizing a custom Brier loss, adapted from the Brier score [17, 18] (see Sec. IV.3) commonly used to evaluate probabilistic forecasts. For a given set of unique observed event times $T=\{T_{j}:\delta_{j}=1\}$, the loss is defined as
$\mathcal{L}=\sum_{t\in T}\sum_{i}\Big[$ $I(T_{i}\leq t\wedge\delta_{i}=1)\,(S(t|\mathbf{x}_{i}))^{2}$ (11)
@@
where $S(t|\mathbf{x}_{i})$ denotes the survival probability predicted for subject $i$, and $I(\cdot)$ is the indicator function.
@@ IV.3 Experimental setup
The procedure consists of two phases. First, for each candidate distribution, we perform 5-fold cross-validation on a training set comprising $80\%$ of the available data, with the remaining $20\%$ held out for testing. Each split preserves the original ratio of observed to censored instances. Cross-validation is used to select the optimal architecture and hyperparameters (see Appendix A), based on average validation performance measured over 100 trials. For hyperparameter search, we use the Optuna framework [19], which implements efficient strategies for exploring the hyperparameter space.
After hyperparameter optimization, models are retrained on the full training set and evaluated on the held-out test set. For each performance metric, mean and standard deviation are estimated via bootstrap resampling: 100 datasets of equal size are drawn with replacement from the test set, each preserving the censoring ratio.
@@
Antolini’s C-index.  We use the time-dependent concordance index (also referred to as Antolini’s C-index) [20] as the performance metric of choice, for both validation and testing. It is a rank statistic that measures agreement between predicted risks and observed survival times as the probability that, among two comparable individuals, the one experiencing the event earlier is assigned a higher predicted risk (or equivalently a lower survival probability). While time-independent in its original formulation [21], Antolini’s formulation generalizes the measure to time-dependent predictions:
@@
where $w(T_{i})$ are inverse probability of censoring weights (IPCW) [22], used to correct for bias introduced by censoring. This time-dependent version is equivalent to the ROC-AUC and is more appropriate for evaluating models whose predictions vary with time.
@@
where lower values indicate better overall predictive accuracy. In our experiments, the IBS is used as an additional metric during the testing phase only.
@@
FLChain:  Data from the Mayo Clinic study of free light chain (FLC) testing [27], including 7 874 patients and 10 features; $72\%$ of cases are censored.
@@
where $x_{i}$ are features. This construction yields non-proportional hazards by design (see Appendix C for more details).
@@ V Results
Figure 2: Performance across clinical datasets. Scatterplots with error bars for C-index (↑, brown, top row) and IBS (↓, dark blue, bottom row) across four datasets (GBSG2, Framingham, SUPPORT2, FLChain). For each dataset, three models are shown: Lévy, inverse Gaussian, and CoxPH. Mean values and standard deviations are computed via 100 bootstrap iterations.
Figure 3: Performance on the NonPH dataset. Scatterplots with error bars showing C-index (↑, brown) and IBS (↓, dark blue) for three models (Lévy, inverse Gaussian, CoxPH). Mean values and standard deviations are computed via 100 bootstrap iterations.
@@ V.1 Model performance
Test results are summarized in Figs. 23. Across datasets and distributions, DeepFHT demonstrates competitive performance, consistently matching or surpassing Cox regression in terms of Antolini’s C-index. Models based on the more physically interpretable Lévy and inverse Gaussian distributions achieve the highest scores on three of the four real-world datasets, while also exhibiting strong robustness and stability at test time.
@@
On the synthetic NonPH dataset, DeepFHT significantly outperforms Cox regression with all tested distributions, underscoring the gains achievable in scenarios with nonlinear effects and violations of the proportional hazards assumption.
@@
Within the clinical datasets, the comparable performance of DeepFHT and Cox regression on GBSG2, Framingham, and FLChain suggests that in these cases the assumptions of linearity and proportional hazards may not be restrictive. This points instead to potential limitations in the expressiveness or complexity of the datasets themselves. By contrast, on the higher-dimensional SUPPORT2 dataset, the inverse Gaussian variant achieves a clear improvement over Cox regression, illustrating the advantage of relaxing linear and proportional hazards assumptions in more complex clinical settings.
V.2 Physics-based interpretability
The parametric nature of DeepFHT, grounded in stochastic process theory, provides a natural and physically meaningful basis for interpreting its predictions. As presented in Section  IV.1, for each individual, the model outputs a vector of parameters of the learned underlying diffusion process, effectively representing the subject as a point in a latent parameter space. This space is inherently interpretable: because the influence of each parameter on the process dynamics is known, we are able to distinguish regions of higher or lower risk and to qualitatively compare the trajectories of individuals in this space. Crucially, such analysis can be performed without explicitly computing survival probabilities.
In our case, the parameterization of Lévy (5) and inverse Gaussian (10) distributions with DeepFHT yields a two-dimensional parameter space. This representation makes interpretation visually intuitive and, in particular, enables the use of Euclidean distance to assess similarity between instances, i.e., whether they exhibit comparable dynamics and event-time probabilities. We show this by visualizing empirical event-time distributions in the parameter space via color gradients and observing that the model consistently assigns previously unseen instances (test set data points) to regions populated by individuals with similar observed event times.
Event times in parameter space.  In the parameter spaces of the Lévy and inverse Gaussian models, we visualize the risk associated with each point by interpolating event times from the training data. Only uncensored instances $\mathbf{p}_{i}\in U=\{\mathbf{p}_{i}\mid\delta_{i}=1\}$ are considered, and interpolation is performed by inverse distance weighting:
$T(\mathbf{p})=\frac{\sum_{\mathbf{p}_{i}\in U}w_{i}(\mathbf{p})\,T_{i}}{\sum_{\mathbf{p}_{i}\in U}w_{i}(\mathbf{p})},\qquad w_{i}(\mathbf{p})=\frac{1}{\sqrt{d(\mathbf{p},\mathbf{p}_{i})}},$

### v3

@@ Abstract
We compare DeepFHT with Cox regression using synthetic and real-world datasets. The method achieves predictive accuracy on par with the state-of-the-art approach, while maintaining a physics-based interpretable parameterization that elucidates the relation between input features and risk. This combination of stochastic process theory and deep learning provides a principled avenue for modeling survival phenomena in complex systems.
@@ I Introduction
Survival analysis is central in many applications across medicine, engineering, economics and finance. It concerns time-to-event modeling: given a process that can generate an event of interest (e.g., death from disease, failure due to wear), the goal is to estimate the probability that an event occurs at any time $t>0$ for an individual described by some input variables (or features, or covariates). Unlike standard regression settings, survival data are characterized by censoring, which means that for some instances, the exact event time is not observed (for example, when individuals remain event-free at the end of the study), and only the last recorded follow-up time is available.
Traditional approaches to survival modeling rely on strong statistical assumptions linking input variables and risk. The Cox proportional hazards (CoxPH) model [1] remains the most widely used and best established method. The proportional hazards assumption implies that the instantaneous risk of event for two individuals differs by a constant factor over time. The CoxPH model is also linear, making it clear how each single input variable affects the outcome, but at the expense of missing interactions between features. In its original form, this relation is modeled through a linear regression on the features, though many extensions have been developed to relax linearity and improve performance in high-dimensional settings [2, 3, 4].
@@
Among these, first hitting time (FHT) models are of particular interest for this work. FHT models assume that event times are distributed as the times of first passage of a latent stochastic process, underlying the event-generating process observed, through an absorbing barrier. This approach overcomes the PH assumption, since FHT distributions for many stochastic processes allow for time-varying hazard ratios between different individuals (e.g., Lévy distribution, Inverse Gaussian distribution).
@@
A branch of survival methods seeks to model the distribution of event times directly, without imposing a parametric process model. Examples include HitBoost [10] and DeepHit [11], which use machine learning and deep neural networks to approximate the event-time density $f(t)$ from data. While these approaches often achieve strong predictive performance, they do so in a purely data-driven manner: the estimated distribution is not tied to parameters of an underlying stochastic process. As a result, the learned models provide limited physical or mechanistic interpretability, and it remains difficult to understand how features influence the latent dynamics that give rise to observed survival outcomes. Conversely, parametric FHT alternatives [12, 13, 14] currently suffer from the opposite drawbacks. These models learn the process parameters on which the distribution of first hitting times depends from data, using simpler regression and machine learning techniques, yielding more physically interpretable results at the expense of model expressiveness.
The present work addresses these complementary limitations by introducing DeepFHT, a framework that unites parametric FHT modeling with deep neural networks. In DeepFHT, a neural network maps input features to the parameters of an underlying stochastic process—such as initial condition, drift, and diffusion—that governs the distribution of first hitting times. To our knowledge, this is the first application of modern deep learning to parametric FHT survival models. The hybrid design combines the representational power of neural networks with the mechanistic grounding of FHT processes, yielding closed-form survival and hazard functions while maintaining a direct physical interpretation of parameters. In this way, DeepFHT provides a principled alternative to black-box survival models, providing predictions that are not only accurate but also anchored in a dynamical description that links covariates to the latent process generating events.
@@
We demonstrate the capability of this framework on synthetic and publicly available clinical survival datasets, comparing predictive power against the CoxPH model. DeepFHT achieves competitive performance across clinical benchmarks and clearly outperforms Cox regression on synthetic data designed to probe the model’s capabilities in scenarios with nonlinearities and non-proportional hazards.
Finally, we assess interpretability by examining the correlations between predicted process parameters, empirical risk, and features. The results show that the similarity between patients outcomes is intuitively encoded as distance in the space of parameters, and that DeepFHT can recover clinically meaningful relationships and has the potential to highlight new correlations in complex, high-dimensional data.
@@
The remainder of the paper is organized as follows. Sections II and III provide theoretical background, reviewing the fundamentals of survival analysis and introducing the relevant aspects of stochastic process theory. Section IV presents the proposed models and the experimental setup, while the results and the physics-based interpretability of the models are discussed in Section V.
@@ II Survival analysis framework
Survival datasets are typically represented as triplets $(\mathbf{x}_{i},\delta_{i},T_{i})$ for $i=1,\dots,n$, where:
  • $n\in\mathbb{N}$ is the dataset size;
  • $\mathbf{x}_{i}\in\mathbb{R}^{m}$ is a vector of $m$ features;
  • $\delta_{i}\in\{0,1\}$ is an indicator that denotes whether the event was observed ($\delta_{i}=1$) or the instance was right censored, i.e., end of observation without occurrence of the event ($\delta_{i}=0$);
  • $T_{i}\in\mathbb{R}_{>0}$ is the time of event occurrence or censoring.
The goal of survival analysis is to predict the probability distribution of event times $f(t)$ and some relevant, related functions:
@@
The survival function describes the probability of surviving beyond time $t$, while the hazard captures the instantaneous risk of event occurrence, given survival up to $t$. Estimation typically focuses on conditional forms $S_{\bm{\theta}}(t|\mathbf{x})$ and $h_{\bm{\theta}}(t|\mathbf{x})$, which quantify the dependence of survival dynamics on the features, through a set of parameters $\bm{\theta}$. If one assumes independent samples and non-informative censoring, the parameters can be determined with a maximum likelihood estimation (MLE) from the data [15].
A widely adopted semi-parametric approach for modeling how covariates affect the time-to-event distribution is the Cox Proportional Hazards model [1]. It assumes that the hazard rate for an individual with features vector $\mathbf{x}$ is proportional to a baseline hazard function $h_{0}(t)$ and can be factorized as:
$h(t\mid\mathbf{x})=h_{0}(t)\exp\left\{\bm{\beta}^{\top}\mathbf{x}\right\},\quad S(t\mid\mathbf{x})=\left[S_{0}(t)\right]^{\exp\left\{\bm{\beta}^{\top}\mathbf{x}\right\}}.$ (1)
This assumption implies that the ratio of hazards for any two individuals is constant over time, depending only on their covariates and not on $t$ itself. Unlike fully parametric methods, CoxPH does not require specifying a functional form for the baseline hazard during the MLE, since the time-dependent components cancel out in the likelihood.
@@ III First hitting time models for survival analysis
A natural probabilistic representation of survival processes is provided by the theory of first hitting times. Consider a one-dimensional continuous-time stochastic process $\{X(t)\}_{t\geq 0}$, with $X(t)\in\mathbb{R}$, described by a probability distribution $p(x,t)$. Let $\Gamma=\{x\leq x_{b}\}\subset\mathbb{R}$ be a target set and $X(t=0)=x_{0}>x_{b}$ the initial condition. In this setup, $x_{b}$ can be considered as a barrier between the target set and the initial condition and this is how it will be referred to in the rest of the paper. The first hitting time is then
@@
that is, the earliest time at which the process crosses the barrier in $x_{b}$. When the barrier is absorbing (i.e., killing the process upon transition), the distribution $f(T)$ of first hitting times coincides with the event time distribution in survival analysis. This condition is easily imposed by the Dirichlet condition $p(x_{b},t\mid x_{0})=0$. Without losing generalizability, we will always consider positive initial conditions $x_{0}>0$ and an absorbing barrier in $x_{b}=0$. This amounts to a generic choice for a process in $X(t)\in\mathbb{R}$.
Since all relevant survival functions can be derived from the distribution of first hitting times $f(T)$, it is crucial to have analytical solutions for this quantity. In this work, we consider two stochastic processes whose theory is well developed and for which an analytical solution exists: the Brownian motion and the arithmetic Brownian motion (Brownian motion with drift). For a more complete and detailed discussion of FHT theory and models, including but not limited to their application to survival analysis, we refer to [16, 12, 17].
@@ III.1 First hitting time and survival distributions for Brownian motion
    the distribution of first hitting times (i.e., event times).
@@ III.2 First hitting time and survival distributions for arithmetic Brownian motion
with a constant diffusion coefficient $D>0$. For an initial condition $p(x,0\mid x_{0})=\delta(x-x_{0})$, the free space solution is Gaussian with the mean shifted linearly in time:
@@
When introducing an absorbing barrier at $x_{b}=0$, it becomes, via the method of images:
@@
    where $\Phi(\cdot)$ denotes the standard normal cdf;
@@
    which corresponds to an inverse Gaussian distribution.
@@ IV.1 DeepFHT
The dimension $d$ of the output vector $\mathbf{p}$ depends on the parametrization choice. As anticipated in Section III, in this paper we will focus on Brownian and arithmetic Brownian motion, identified by their FHT distributions 6,11, with the following parameterizations:
@@
Notably, both parameterizations above allow for non-proportional hazards, reflecting the flexibility of FHT-based models to capture complex time dependencies in risk. Figure 1 illustrates a representative model output, where survival functions are computed from the estimated distributional parameters for a set of individuals. It is also worth noting that this model can easily be extended to any other probability distribution of choice. It is sufficient to specify the functional form and its parameterization and to modify the size of the output layer accordingly.
Physics-based interpretability
A distinctive feature of DeepFHT is its physics-based interpretability. Because the model is parametrized by quantities with direct physical meaning (initial condition, drift, and diffusion) each individual is represented as a point in a latent parameter space whose geometry reflects the underlying process dynamics. The influence of every parameter on event-time distribution is analytically known, allowing regions of higher or lower risk to be distinguished and trajectories of individuals to be compared qualitatively in this space. In the specific Lévy (6) and inverse Gaussian (11) cases, the resulting two-dimensional parameterization makes these relationships visually intuitive and enables quantitative notions of similarity (e.g., Euclidean distance) between subjects.
To visualize risk in parameter space, we interpolate event times from the set of uncensored training instances $U=\{\mathbf{p}_{i}\mid\delta_{i}=1\}$ via inverse-distance weighting:
$T(\mathbf{p})=\frac{\sum_{\mathbf{p}_{i}\in U}w_{i}(\mathbf{p})\,T_{i}}{\sum_{\mathbf{p}_{i}\in U}w_{i}(\mathbf{p})},\qquad w_{i}(\mathbf{p})=\frac{1}{\sqrt{d(\mathbf{p},\mathbf{p}_{i})}}$ (12)
Here $d(\cdot,\cdot)$ is a distance in the model parameter space. The resulting $T(\mathbf{p})$ acts as a proxy for risk and supports distance-based similarity analyses.
@@ IV.2 Loss function: Brier loss
Training is performed by minimizing a custom Brier loss, adapted from the Brier score [18, 19] (see Sec. IV.3) commonly used to evaluate probabilistic forecasts. The loss is defined as
$\mathcal{L}=\sum_{t\in U}\sum_{i}\Big[$ $I(T_{i}\leq t\wedge\delta_{i}=1)\,(S(t|\mathbf{x}_{i}))^{2}$ (13)
@@
where $U=\{T_{j}:\delta_{j}=1\}$ is the set of unique event times where predictions are evaluated. $S(t|\mathbf{x}_{i})$ denotes the survival probability predicted for subject $i$, and $I(\cdot)$ is the indicator function.
@@ IV.3 Experimental setup
The procedure consists of two phases. First, for each candidate distribution, we perform 5-fold cross-validation on a training set comprising $80\%$ of the available data, with the remaining $20\%$ held out for testing. Each split preserves the original ratio of observed to censored instances. Cross-validation is used to select the optimal architecture and hyperparameters (see Appendix A), based on average validation performance measured over 100 trials. For hyperparameter search, we use the Optuna framework [20], which implements efficient search strategy for exploring the hyperparameter space.
After hyperparameter optimization, models are retrained on the full training set and evaluated on the held-out test set. For each performance metric, mean and standard deviation are estimated via bootstrap with resampling: 100 datasets of equal size are drawn with replacement from the test set, each preserving the censoring ratio.
@@
Antolini’s C-index.  We use the time-dependent concordance index (also referred to as Antolini’s C-index) [21] as the performance metric of choice, for both validation and testing. It is a rank statistic that measures agreement between predicted risks and observed survival times as the probability that, among two comparable individuals, the one experiencing the event earlier is assigned a higher predicted risk (or equivalently a lower survival probability). Antolini’s C-index generalizes the original Harrel’s C-index [22] to models where the predicted risk ranking between individuals is not fixed (like in CoxPH) but changes at different times.
@@
where $w(T_{i})$ are inverse probability of censoring weights (IPCW) [23], used to correct for bias introduced by censoring. This time-dependent version is a censoring-aware equivalent to the ROC-AUC metrics and it is more appropriate for evaluating models whose predictions vary with time.
@@
where lower values indicate better overall predictive accuracy and a Brier score of 0.25 corresponds to a completely uninformative model. In our experiments, the IBS is used as an additional metric during the testing phase only.
@@
where $x_{i}$ are features. This construction yields non-proportional hazards by design (see Appendix B for more details).
@@ V Results
Figure 2: Performance across clinical and synthetic datasets. Scatterplots with error bars for C-index (↑, brown, top row) and IBS (↓, dark blue, bottom row) across datasets (GBSG2, Framingham, SUPPORT2, NonPH). For each dataset, three models are shown: Lévy, inverse Gaussian, and CoxPH. Mean values and standard deviations are computed via 100 bootstrap iterations.
@@ V.1 Model performance
Predictive test results are summarized in Fig. 2. Across datasets, models based on the more physically interpretable Lévy and inverse Gaussian distributions demonstrates competitive performance, consistently matching or surpassing Cox regression in terms of Antolini’s C-index, exhibiting strong robustness and stability at test time.
@@
On the synthetic NonPH dataset, DeepFHT significantly outperforms CoxPH with all tested distributions, underscoring the gains achievable in scenarios with nonlinear effects and violations of the proportional hazards assumption.
@@
Within the clinical datasets, the comparable performance of DeepFHT and CoxPH on GBSG2 and Framingham suggests that in these cases the assumptions of linearity and proportional hazards may not be too restrictive. This points instead to potential limitations in the expressiveness or complexity of the datasets themselves. By contrast, on the higher-dimensional SUPPORT2 dataset, the inverse Gaussian variant achieves a clear improvement over Cox regression, illustrating the advantage of relaxing linear and proportional hazards assumptions in more complex settings.
V.2 Empirical analysis of parameter-space interpretability
Building upon the interpretability introduced in Section IV.1, we now examine how the learned parameters encode survival dynamics in practice. Specifically, we visualize event times and clinical features in the two-dimensional parameter spaces of the Lévy and inverse Gaussian models.